Maria’s age is 20. Her father’s age is 50. In how many years will Maria's father be twice as old as Maria will be?

Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52794)

.

Let x be "in how many years . . . "


Then you have this equation to find x


    2*(20+x) = 50+x.


Simplify it step by step


    40 + 2x = 50 + x

    2x - x = 50 - 40

     x     = 10.


ANSWER.  In 10 years.

Solved.

-----------------

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problems
    - Selected age word problems from the archive
    - Age problems for mental solution
    - Miscellaneous age problems
    - OVERVIEW of lessons on age problems
in this site.

Read them and become an expert in solving age problems.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.


Answer by greenestamps(13200)



An informal solution, using logical reasoning....

The difference in their ages is 50-20 = 30 years.
So her father will be twice as old as she is when she is 30 and he is 60.
Since they are now 20 and 50, that will be 10 years from now.